OpenStudy (anonymous):

Calculate the average rate of change for the function f(x) = x4 + 3x3 − 5x2 + 2x − 2, from x = −1 to x = 1.

OpenStudy (anonymous):

@Marcorie

OpenStudy (anonymous):

what are you answers? Cause i already have mine

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

no. And what are your options

OpenStudy (anonymous):

its positive 5

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

no f x=1, f(1)= (1)^4 + 3(1)^3 + 5(1)^2 + 2(1) - 2 =

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Last one:)

OpenStudy (anonymous):

k wait one sec

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

(x)=x^3+x^2-2x+1 is this how the problem should look like?

OpenStudy (anonymous):

Yes. That's what I thought

OpenStudy (anonymous):

up for another?

OpenStudy (anonymous):

f(x)=(x+1)^3+ (x+1)^2-2(x+1)+1

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

just wondering is this a test or a quiz?

OpenStudy (anonymous):

Homework :(

OpenStudy (anonymous):

I have summer school

OpenStudy (anonymous):

oh ok well im not going to give you the answer but i will give you this hint, use a graphing calculator to view the quartic function and see the function increasing from x = -2 to -1; decreasing from 0 to 1; increasing from 1 to 2; and increasing from 2 to 4. okay tell me when you find the answer.

OpenStudy (anonymous):

Ok so i got x=0, x=1

OpenStudy (anonymous):

Also, I think the last question i posted was the wrong answer. Bc when I did it I got 1^4+3(1)^3-5(1)^2+2(1)-2= -5

OpenStudy (anonymous):

well thats what i got

OpenStudy (anonymous):

ok. thanks:)